GIven that log36=m, and log65=n express log310 in terms of m and n


1
Expert's answer
2021-08-30T19:06:32-0400

Use the formula "log_a (bc)=log_a b +log_a c":


"log_3 10 =log_3 (2\cdot 5)=log_3 2+log_3 5"

Express "log_3 2" from the first equality:


"log_3 6 = m""log_3 (2\cdot3)=m""log_3 2+log_3 3=m""log_3 2+1=m""log_3 2=m-1"

Express "log_3 5" by using formulas "log_a b=\frac{1}{log_b a}, \frac{log_a b}{log_a c}=log_c b"


"log_3 5=\frac{log_6 5}{log_6 3}=log_6 5\cdot log_3 6=nm"

Thus,


"log_3 10 =log_3 2+log_3 5 =m-1+nm=mn+m-1"

Answer: "mn+m-1"